# Hybrid quantum lattice model: Polaritons, photons, and spin waves propagation

### v2

@@ Appendix A Polaritonic basis
The eigenstates of the Jaynes-Cummings model define the upper (+)(+)( + ) and lower ()(-)( - ) polaritonic basis, which is used to write the Hamiltonian of our model to analyze the effective couplings to propagate the different excitations. A compact way to write the basis is to use the following notation
@@
$\begin{split}\mathbf{I}\hskip 2.84544pt\hat{\sigma}^{+}\mathbf{I}&=\sum_{n,m=0}^{\infty}\sum_{\alpha,\beta=\pm}P^{\dagger(n,\alpha)}P^{(n,\alpha)}\hat{\sigma}^{+}P^{\dagger(m,\beta)}P^{(m,\beta)}\,,\\ &=\sum_{n,m=0}^{\infty}\sum_{\alpha,\beta=\pm}\langle n,\alpha|\hat{\sigma}^{+}|m,\beta\rangle P^{\dagger(n,\alpha)}P^{(m,\beta)}\,,\end{split}$ (28)
@@
Here, the Kronecker delta is chosen such that $m=n-1$
@@
$\begin{split}\hat{\sigma}_{j}^{+}\hat{\sigma}_{j+1}^{-}=\sum_{n,m=1}^{\infty}\sum_{\begin{subarray}{c}\alpha,\alpha^{\prime}=\pm\\ \beta,\beta^{\prime}=\pm\end{subarray}}\big{(}&\rho_{n\alpha}\gamma_{n-1\beta}\rho_{m\alpha^{\prime}}\gamma_{m-1\beta^{\prime}}P_{j}^{\dagger(n,\alpha)}\\ &P_{j}^{(n-1,\beta)}P_{j+1}^{\dagger(m-1,\beta^{\prime})}P_{j+1}^{(m,\alpha^{\prime})}\big{)}\,,\end{split}$
@@
$\begin{split}U_{j}(t)&=\exp\bigg{(}{-\frac{i}{\hbar}\sum_{n=1}^{\infty}\sum_{\alpha=\pm}F_{n}^{\alpha}(t)P_{j}^{\dagger(n,\alpha)}P_{j}^{(n,\alpha)}}\bigg{)}\,,\\ &={\sum_{n=1}^{\infty}\sum_{\alpha=\pm}\mathrm{e}^{-\frac{i}{\hbar}F_{n}^{\alpha}(t)}P_{j}^{\dagger(n,\alpha)}P_{j}^{(n,\alpha)}}\,,\end{split}$
@@ Appendix B Truncation error
In the TEBD algorithm, a two-site wave function’s split and truncate procedure in a mixed canonical form involves the singular value decomposition (SDV) of a 4-rank multidimensional array. The SVD allows us to compute ${\rm norm_{old}}=\sqrt{\sum_{a}\lambda_{a}^{2}}$ where $\lambda_{a}$ are the Schmidt coefficients. Then, we proceed with the truncation by keeping the Schmidt coefficients larger than a given threshold $\epsilon_{0}$, which we fix to $\epsilon_{0}=10^{-6}$ in our algorithm. The new bond dimension of the updated matrix product state (MPS) is computed as ${\rm min}(\chi_{\rm max},N_{\rm kept})$ where $N_{\rm kept}$ is the number of kept singular values, and $\chi_{\max}$ is the maximum bond dimension of the MPS which we fix to $\chi_{\rm max}=4$. Introducing the threshold value $\epsilon_{0}$ allows an adaptive bond dimension along the quantum lattice, thus saving computational time. Once the truncation has been done, we compute ${\rm norm_{new}}=\sqrt{\sum_{a=1}^{N_{\rm kept}}\lambda_{a}^{2}}$ and $1-2\epsilon_{n}$, where $\epsilon_{n}=1-({\rm norm_{new}}/{\rm norm_{old}})^{2}$. The described procedure is done each time a two-body unitary is applied following the Suzuki-Trotter decomposition [64] of the evolution operator, and the quantity $1-2\epsilon_{n}$ is saved in a vector. The total truncation error is computed as $\epsilon_{\rm trunc}=1-\prod_{n}(1-2\epsilon_{n})$ [65].

### v3

@@ Appendix A Polaritonic basis
The eigenstates of the Jaynes-Cummings model define the upper $(+)$ and lower ()(-) polaritonic basis, which is used to write the Hamiltonian of our model to analyze the effective couplings to propagate the different excitations. A compact way to write the basis is to use the following notation
@@
[…]
@@
Here, the Kronecker delta is chosen such that m=n1m=n-1
@@
[…]
@@
where $F_{n}^{\alpha}(t)=E_{n}^{\alpha}t$. As a result of the previous, we only need to compute two terms: $\textrm{I})\hskip 2.84544ptU_{j}^{\dagger}(t)P_{j}^{\dagger(n,\alpha)}P_{j}^{(n-1,\beta)}U_{j}(t)$ and $\textrm{II)}\hskip 2.84544ptU_{j+1}^{\dagger}(t)P_{j+1}^{\dagger(m-1,\beta^{\prime})}P_{j+1}^{(m,\alpha^{\prime})}U_{j+1}(t)$. Without loss of generality, we calculate the terms I and II for a general index site $j=l$, and we highlight that the polaritonic branch indexes and the number of polaritons are independent for every operator of these terms. Consequently, for the local evolution operators, we introduce the indexes $\eta$ and $\mu$ for the number of excitations and $\zeta$ and $\xi$ for the polaritonic branches.
@@ Appendix B Truncation error
In the TEBD algorithm, a two-site wave function’s split and truncate procedure in a mixed canonical form involves the singular value decomposition (SDV) of a 4-rank multidimensional array. The SVD allows us to compute ${\rm norm_{old}}=\sqrt{\sum_{a}\lambda_{a}^{2}}$ where $\lambda_{a}$ are the Schmidt coefficients. Then, we proceed with the truncation by keeping the Schmidt coefficients larger than a given threshold $\epsilon_{0}$, which we fix to $\epsilon_{0}=10^{-6}$ in our algorithm. The new bond dimension of the updated matrix product state (MPS) is computed as ${\rm min}(\chi_{\rm max},N_{\rm kept})$, where $N_{\rm kept}$ is the number of kept singular values, and $\chi_{\max}$ is the maximum bond dimension of the MPS, which we fix to $\chi_{\rm max}=4$ for single excitation propagation and $\chi_{\rm max}=50$ for the excitation swap. Introducing the threshold value $\epsilon_{0}$ allows an adaptive bond dimension along the quantum lattice, thus saving computational time. Once the truncation has been done, we compute ${\rm norm_{new}}=\sqrt{\sum_{a=1}^{N_{\rm kept}}\lambda_{a}^{2}}$ and $1-2\epsilon_{n}$, where $\epsilon_{n}=1-({\rm norm_{new}}/{\rm norm_{old}})^{2}$. The described procedure is done each time a two-body unitary is applied following the Suzuki-Trotter decomposition [75] of the evolution operator, and the quantity $1-2\epsilon_{n}$ is saved in a vector. The total truncation error is computed as $\epsilon_{\rm trunc}=1-\prod_{n}(1-2\epsilon_{n})$ [80].
@@ Appendix C Polaritonic branch populations
Appendix D Swap sensitivity
Figure D1: Swap parameters sensitivity, polariton $\rightarrow$ atomic-wave: Average excitation number of the $j$th TLS and the average photon number as functions of time for the left section ($\ell$), shown in panels (a) and (b), and for the right section ($r$), shown in panels (c) and (d), respectively. (e) Average polariton number as a function of time and the site index $j$ (with $2\leq j\leq 31$). The parameters in the left section are $\omega_{\ell}=\omega^{\ell}_{0}=1$, $\Delta_{\ell}=\omega^{\ell}_{0}-\omega_{\ell}=0$, $v_{\ell}=0.05\,\omega_{\ell}$, $g_{\ell}=4v_{\ell}$, $\omega_{\rm{A}}=\omega^{\ell}_{pol}=\omega_{\ell}-g_{\ell}$, $\lambda=-v_{\ell}/\sqrt{2}$, $\lambda^{\prime}_{\rm{C}}=-\,v_{\ell}/2\sqrt{2}$. In the right section, the parameter are $\omega^{r}_{0}=\omega^{\ell}_{pol}$, $\omega_{r}=50\,\omega^{r}_{0}$, $\Delta_{r}=\omega^{r}_{0}-\omega_{r}$, $g_{r}=g_{\ell}$, $\theta_{1}=\arctan{(2g_{r}/\Delta_{r})}$, $v_{r}=v_{\ell}/2$. Additional parameters of the numerical simulation are $\tau=10^{-3}v_{\ell}^{-1}$, $n_{\rm{max}}=2$, $\chi_{\rm max}=50$, and $L=31$.
We present the swap dynamics when the matching and resonance conditions are detuned from their optimal values [see Eqs.(22) and (25) in the main text]. Figure D1 shows the excitation dynamics when the impedance-matching condition is modified to $\lambda^{\prime}_{\rm{C}}=\lambda_{\rm{C}}/2$, while all remaining parameters are identical to those used in Fig. 6. In this case, the conversion from a polariton to an atomic wave still occurs, but the process is no longer ideal: a partial reflection appears on the left side of the interface. Moreover, in the right section of the lattice, the atomic wave propagates with a reduced group velocity, as indicated by the smaller slope of the wavefront in Fig. D1(e).
Figure D2 highlights the sensitivity of the swapping mechanism to the resonance condition, which in the optimal case requires the polariton frequency on the left side to match the transition frequency of the TLSs on the right, i.e., $\omega^{r}_{0}=\omega^{\ell}_{pol}$. In Fig. D2, we detune this condition by setting $\omega^{r}_{0}=\omega^{\ell}_{pol}/2$, while keeping all other parameters equal to those of Fig. 6. Under this detuned resonance, the incoming polariton wave is completely reflected, and the atomic excitation fails to propagate into the right-hand lattice. This behavior is consistent with Figs. D2(c)–D2(e), which show the absence of excitation transfer across the interface.
Figure D2: Swap parameters sensitivity, polariton $\rightarrow$ atomic-wave: Average excitation number of the $j$th TLS and the average photon number as functions of time for the left section ($\ell$), shown in panels (a) and (b), and for the right section ($r$), shown in panels (c) and (d), respectively. (e) Average polariton number as a function of time and the site index $j$ (with $2\leq j\leq 31$). The parameters in the left section are $\omega_{\ell}=\omega^{\ell}_{0}=1$, $\Delta_{\ell}=\omega^{\ell}_{0}-\omega_{\ell}=0$, $v_{\ell}=0.05\,\omega_{\ell}$, $g_{\ell}=4v_{\ell}$, $\omega_{\rm{A}}=\omega^{\ell}_{pol}=\omega_{\ell}-g_{\ell}$, $\lambda=\lambda_{\rm{C}}=-v_{\ell}/\sqrt{2}$. In the right section, the parameter are $\omega^{r}_{0}=\omega^{\ell}_{pol}/2$, $\omega_{r}=50\,\omega^{r}_{0}$, $\Delta_{r}=\omega^{r}_{0}-\omega_{r}$, $g_{r}=g_{\ell}$, $\theta_{1}=\arctan{(2g_{r}/\Delta_{r})}$, $v_{r}=v_{\ell}/2$. Additional parameters of the numerical simulation are $\tau=10^{-3}v_{\ell}^{-1}$, $n_{\rm{max}}=2$, $\chi_{\rm max}=50$, and $L=31$.